# Quantifying Calibration Error in Modern Neural Networks through Evidence Based Theory

### v1

@@ 7.3 Discussion
The behavior of the metrics for the CIFAR-10 model can be explained by the fact that the model requires around 10 epochs to be well-trained. After this point, as shown in Figure 3(b), the model begins to overfit, as indicated by the increasing gap between the training and validation losses. This overfitting leads to poor generalization and explains why the metrics—trust, disbelief, and uncertainty—are optimal for both pre- and post-calibration models during the early epochs, but deteriorate sharply for the pre-calibration model after about 10 epochs. Specifically, trust decreases, and disbelief and uncertainty increase for the pre-calibration model as overfitting sets in. In contrast, the post-calibration model experiences a slight decline in performance but stabilizes quickly, maintaining higher trustworthiness throughout the training.
@@
  • Overfitting in CIFAR-10: The overfitting observed in the CIFAR-10 model results in a more overconfident network. This overconfidence leads to poor probability predictions, which causes trust to decrease and disbelief to increase, as evidenced by the worsening of these metrics in the pre-calibration model. After around 20 epochs, trust in the pre-calibration model drops to around 0.48, and disbelief peaks at 0.58, remaining relatively constant thereafter (as shown in the results).
@@
  • Post-Calibration Stability: Temperature scaling mitigates the effects of overfitting by adjusting the model’s confidence levels. This explains why the post-calibration model remains stable after the initial performance dip and avoids the severe deterioration seen in the pre-calibration model. After calibration, trust decreases only slightly and stabilizes around 0.7 after 20 epochs, while disbelief and uncertainty also stabilize at much lower levels compared to the pre-calibration model.
For the MNIST model, the dynamics are different. As shown in Figure 3(a), both training and validation losses remain low and close to each other throughout the training process, indicating little to no overfitting. This explains why we do not observe the same pattern of improvement followed by deterioration as seen in CIFAR-10.
@@
  • MNIST Training Behavior: The MNIST model reaches its best performance relatively early in training, and there is little difference in behavior between the pre- and post-calibration models. The absence of significant overfitting means that trust remains relatively constant throughout training for both models, and disbelief and uncertainty are minimal. The difference between the pre- and post-calibration models is less pronounced because the MNIST task is simpler and requires fewer epochs to reach optimal performance.
@@
  • Impact of Overfitting on Trustworthiness: Overfitting leads to overconfident neural networks, which in turn produce poorly calibrated probability predictions. This decreases belief (or trust) in the network’s predictions and increases disbelief and uncertainty. In the CIFAR-10 model, these effects are evident, with trust decreasing and disbelief increasing as overfitting sets in. However, temperature scaling corrects for this by recalibrating the model’s confidence, which helps the post-calibration model remain trustworthy even after the base model begins to overfit.
In conclusion, the key to understanding the results is the role of overfitting in the CIFAR-10 model, which leads to a steep decline in performance for the pre-calibration model. Temperature scaling helps mitigate this, resulting in a more stable and trustworthy model. In contrast, the MNIST model does not experience significant overfitting, and thus the difference between pre- and post-calibration models is less marked.
@@ 8 Conclusion
The evaluation of our framework demonstrated its effectiveness in maintaining stable trustworthiness over time, particularly when applied to calibrated models. Our experimental results on MNIST and CIFAR-10 datasets highlight the improvements in trust, disbelief, and uncertainty metrics after calibration, showcasing the utility of temperature scaling in enhancing model reliability. By dynamically updating trustworthiness during inference, this framework has the potential to contribute to the ethical deployment of AI systems in critical domains such as healthcare and autonomous systems.
Future Works. To further refine our approach, future work will focus on elaborating the scheme to make uncertainty solely dependent on the amount of data. This would ensure that uncertainty reflects the quantity of evidence available, preventing redundancy with disbelief. Additionally, we aim to enhance the dynamic aspects of the framework, enabling real-time trustworthiness updates during the operational phase. Another key area for exploration is understanding the impact of the number of clusters on the results, as more clusters allow for greater dynamism but reduce the amount of evidence available for quantification for each cluster.

### v2

@@ VII-C Discussion
The behavior of the metrics for the CIFAR-10 model can be explained by the fact that the model requires around 10 epochs to be well-trained. After this point, as shown in Figure 3(b), the model begins to overfit, as indicated by the increasing gap between the training and validation losses. This overfitting leads to poor generalization and explains why the metrics—trust, distrust, and uncertainty—are optimal for both pre- and post-calibration models during the early epochs, but deteriorate sharply for the pre-calibration model after about 10 epochs. Specifically, trust decreases, and distrust and uncertainty increase for the pre-calibration model as overfitting sets in. In contrast, the post-calibration model experiences a slight decline in performance but stabilizes quickly, maintaining higher trustworthiness throughout the training.
Overfitting in CIFAR-10
@@
The overfitting observed in the CIFAR-10 model results in a more overconfident network. This overconfidence leads to poor probability predictions, which causes trust to decrease and distrust to increase, as evidenced by the worsening of these metrics in the pre-calibration model. After around 20 epochs, trust in the pre-calibration model drops to around 0.48, and distrust peaks at 0.58, remaining relatively constant thereafter (as shown in the results).
Post-Calibration Stability
@@
Temperature scaling mitigates the effects of overfitting by adjusting the model’s confidence levels. This explains why the post-calibration model remains stable after the initial performance dip and avoids the severe deterioration seen in the pre-calibration model. After calibration, trust decreases only slightly and stabilizes around 0.7 after 20 epochs, while distrust and uncertainty also stabilize at much lower levels compared to the pre-calibration model.
For the MNIST model, the dynamics exhibit similarities to those observed in CIFAR-10 but with a less pronounced effect. As shown in Figure 3(a), the gap between training and validation losses increases throughout training, though not as significantly as in CIFAR-10. Toward the later epochs, the gap appears to stabilize, suggesting a lower degree of overfitting compared to CIFAR-10. This explains why we do not observe the same sharp pattern of improvement followed by deterioration seen in CIFAR-10, where overfitting is more severe.
MNIST Training Behavior
@@
The MNIST model reaches its best performance relatively early in training, and there is little difference in behavior between the pre- and post-calibration models. The absence of significant overfitting means that trust remains relatively constant throughout training for both models, and distrust and uncertainty are minimal. The difference between the pre- and post-calibration models is less pronounced because the MNIST task is simpler and requires fewer epochs to reach optimal performance.
Impact of Overfitting on Trustworthiness
@@
Overfitting leads to overconfident neural networks, which in turn produce poorly calibrated probability predictions. This decreases belief (or trust) in the network’s predictions and increases distrust and uncertainty. In the CIFAR-10 model, these effects are evident, with trust decreasing and distrust increasing as overfitting sets in. However, temperature scaling corrects for this by recalibrating the model’s confidence, which helps the post-calibration model remain trustworthy even after the base model begins to overfit.
@@ VIII Conclusion
The evaluation of our framework demonstrated its effectiveness in maintaining stable trustworthiness over time, particularly when applied to calibrated models. Our experimental results on MNIST and CIFAR-10 datasets highlight the improvements in trust, distrust, and uncertainty metrics after calibration, showcasing the utility of temperature scaling in enhancing model reliability. By dynamically updating trustworthiness during inference, this framework has the potential to contribute to the ethical deployment of AI systems in critical domains such as healthcare and autonomous systems.
Future Works. Future work will focus on elaborating the scheme to make uncertainty solely dependent on the amount of data. This would ensure that uncertainty reflects the quantity of evidence available, preventing redundancy with disbelief. Additionally, we aim to enhance the dynamic aspects of the framework, enabling real-time trustworthiness updates during the operational phase. Another key area for exploration is understanding the impact of the number of clusters on the results, as more clusters allow for greater dynamism but reduce the amount of evidence available for quantification for each cluster.